OpenStudy (anonymous):

Write a system of equations for the following problems and solve: The receipts from 300 tickets for a high school program were $650. Adults were charged $2.50 each and students $1.50. How many tickets of each kind were sold?

OpenStudy (anonymous):

Let's start off by writing the equations. Can you identify what the two equations would need to be about to start off?

OpenStudy (anonymous):

Um lets see I am not sure..

OpenStudy (texaschic101):

a = adult tickets s = student tickets a + s = 300 ---> s = 300 - a 2.5a + 1.5s = 650 2.5a + 1.5(300 - a) = 650 2.5a + 450 - 1.5a = 650 2.5a - 1.5a = 650 - 450 a = 200 a + s = 300 200 + s = 300 s = 300 - 200 s = 100 check.. 2.5a + 1.5s = 650 2.5(200) + 1.5(100) = 650 500 + 150 = 650 650 = 650 (correct) adult tickets (a) = 200 student tickets (s) = 100

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Thank you so much perfect explanation

OpenStudy (texaschic101):

glad to help...sorry it took so long...at first, I couldn't get it to equal, but I finally caught my mistake and all is good

OpenStudy (anonymous):

that is ok! Thank you. :)